OpenStudy (anonymous):

The sum of n terms of an arithmetic sequence with first term 3 and common difference 4 is 1830. Calculate the value of n

OpenStudy (anonymous):

My working out: S1830 =n/2 (2x3+(n-1)x4) 1830 = 2n (6+ n x4) n = 1830/2 x 24

OpenStudy (anonymous):

but i got answer completely wrong lol

OpenStudy (anonymous):

re-check your second step..

OpenStudy (anonymous):

\[Sn = \frac{ n }{ 2 } \times ( 2a + (n-1)d)\] \[1830 = \frac{ n }{ 2 } \times ( 2*3 + (n-1) 4)\]

OpenStudy (anonymous):

n/2(2+4n) n+2n^2 =1830 2n^2+n-1830=0 (2n+61)(n-30)=0

OpenStudy (anonymous):

\[1830*2 = n(6 +4n -4)\] \[3660 = n (4n +2)\] \[3660 = 4n ^{2} + 2n\] \[2n ^{2} + n - 1830 = 0 \]